OpenStudy (anonymous):

-3 1/5 divided by 2/15

OpenStudy (gorv):

\[-3\frac{ 1 }{ 5 }=\frac{ -15+1 }{ 5 }=\frac{ -14 }{ 5 }\]

OpenStudy (anonymous):

this are the choices 1/24 -1/24 24 -24

OpenStudy (gorv):

loll that it should be -(3 1/5) = -16/5

OpenStudy (anonymous):

-4.5 is the answer

OpenStudy (gorv):

\[\frac{ -16 }{ 5 }\div \frac{ 2 }{15}=\frac{ -16 }{ 5 }*\frac{ 15 }{ 2 }=-24\]

OpenStudy (anonymous):

thnx guys

OpenStudy (gorv):

welcome